OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 3.0 centimeters/minute while the area of the triangle is increasing at a rate of 2.0 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 11.5 centimeters and the area is 91.0 square centimeters?

OpenStudy (anonymous):

\(\large A=\frac{1}{2}bh \rightarrow A'=\frac{1}{2}(b'h+bh')\) since h=11.5 and A=91, you can find b. then you can replace all those other given values to find b' (the rate at which the base is changing)...

OpenStudy (anonymous):

i tried but i kept getting the wrong answer

OpenStudy (anonymous):

what did you get for b ???

OpenStudy (anonymous):

15.82

OpenStudy (anonymous):

the base is decreasing at 3.78 cm/min is incorrect?

OpenStudy (anonymous):

its coming up as incorrect

OpenStudy (anonymous):

ok... lesseee....

OpenStudy (anonymous):

\(\large A'=\frac{1}{2}(b'h+bh') \rightarrow b'=\frac{2A'-bh'}{h} \)

OpenStudy (anonymous):

91 = (1/2)*(b)*(11.5) so b = 15.826 now... A' = 2 b = 15.826 h' = 3 h = 11.5

OpenStudy (anonymous):

hmmm... i still got b' = -3.781 cm/min...

OpenStudy (anonymous):

ohh i missed the neg sign

OpenStudy (anonymous):

it worked out now thanks!

OpenStudy (anonymous):

yw... btw... my first answer was positive because I said the base is "decreasing at 3.78 cm/min" this is also correct.

OpenStudy (anonymous):

yeah now that i reread it i understood

OpenStudy (anonymous):

great.... :)

OpenStudy (anonymous):

and that is something to watch out for with these related rates problems.....

OpenStudy (anonymous):

im posting another problem that i am having trouble with, do you mind checking it out to see if you can help

OpenStudy (anonymous):

yeah... no prob....